Rick always likes to consume a cup of tea with a serving of noodles at any given time of the day. He earns a monthly allowance of 150. A cup of tea cost him Rs5 and serving of noodles cost Rs 10. Given this what will be Rick’s optimal consumption bundle? (1.5 marks) 2) Next, the price of a serving oof noodles has fallen to Rs9 while the price of tea has gone up to Rs6. How will it impact the choice of Rick? (1.5 marks) 3) Finally, consider instead, that the price of Noodles has fallen to Rs8 while the price of a cup of tea is constant at Rs 5, how will the choice be impacted in this case? Discuss the income and substitution effect in this case. (2 marks) Use diagrams to support your answers


1
Expert's answer
2021-10-12T10:03:27-0400

1.Rick's preference can be expressed as U"=" xy

Monthly income"=" 150(U)

Tea cost Rs.5 (x)

Noodles cost Rs.10(y)

We will use Budget line where by

"I=(P_x)x+(P_y)y"

"150=5x+10y"

Then we will use indifference curve and budget line as tangent

Slope of IC"=" "MU_x =y" and "MU_y=x"

So "\frac{-5}{10}= \frac{-1}{2}"

Slope of IC"=" slope of BL

"\frac {-y}{x}=\frac{-1}{2}"

"2y=x"

We solve for two equations with 2 unknowns

150"=5x+10y" and 2y"=x"

150"=" "5(2y)+10y"

150"=" 20"y"

7.5"=" "y"

For "x:"

2(7.5)"=" "x"

"15=x"


2.Slope of IC"=" slope of BL

"\frac{-y}{x}=\frac{-2}{3}"

1.5y"=x"

"150=6x+9y"

"150=6(1.5y)+9y"

"150=18y"

"y=" "8.3"

For x :

"1.5y=x"

"1.5(8.3)=x"

"x=12.45"

3.Based on income effect and substitution effect Rick may consume more noodles because of it's low price.